lizzie777:

Braelyn is writing an equation to represent the perimeter of a rectangle with the width equal to half the length. She knows that the perimeter is equal to 18 units and needs to find the length and width of the rectangle. Which equations represent this relationship? Check all that apply. x + one-half x = 18 2 x + x = 18 2 (x) + 2 (2 x) = 18 x + one-half x + x + one-half x = 18 2 (x + one-half x) = 18 2 x + one-half x = 18

darkknight:

Which one do you think it is?

darkknight:

Well, from what we know, the perimeter is 18

lizzie777:

I think its is the 2,3

darkknight:

and the width is 1/2(length)

darkknight:

perimeter = 2width + 2length

darkknight:

So try setting it up like this: 2(width)+2(length) = 18 2(1/2 length)+ 2(length)

darkknight:

Do you see where this comes from?

darkknight:

We can write the question is terms of width AND length. Rn we are doing in terms of length

lizzie777:

yes

darkknight:

Okay, so we have 2(1/2 length)+ 2(length) = 18 One variable, so we can solve for that 2(1/2) = 1 so length + 2 length = 3 length length + 2 length = 18

darkknight:

You can replace length for x, does this match any of your answer choices?

lizzie777:

one of them

darkknight:

Yes the 2nd one

darkknight:

Now we can eliminate some answer choices that aren't possible

darkknight:

we have stated that the length is x, and that is the only variable. So we know 2 x + x = 18 is true. So can this be true then? x + one-half x = 18 ???

darkknight:

@lizzie777 What I am asking is that knowing that the second answer choice is right. Can the first also be right?

lizzie777:

I could be

darkknight:

x + one-half x = 18 2 x + x = 18 Well since both of the above equations are equal to 18, you can set them equal to each other. x+ 1/2x = 2x+x add like terms 3x = 3/2x

darkknight:

3x = 3/2 x only when x is zero, right? but the perimeter is 18 so answer 1 is incorrect

darkknight:

Lets make it quicker. You want equations that are similar to 2 x + x = 18, can you spot any off the bat?

lizzie777:

its the second one

darkknight:

there can be more answer choices than the second one

lizzie777:

the 4 one?

darkknight:

I am asking if any of the other answer choices are like the second one that is, rewritting 2 x + x = 18

darkknight:

Yes the fourth one is also correct

darkknight:

Can you spot anymore? hint: try looking at the 5th one

lizzie777:

I know its not the first one

lizzie777:

the 5 one look like it could be

darkknight:

try rewritting the 5th one. It is also equal to 18. So set the 5th one equal to the 2nd one and see if you can get something like 1 =1

lizzie777:

I have to go I''ll be right back

darkknight:

here, Ill help you. 2 (x + one-half x) = 18 2x+x=18 2(x+1/2x) = 2x+x 2x+x = 2x+x 0 = 0

darkknight:

So that means that the 5th one is also correct

darkknight:

You are done! Good job persisting. correct answers: 2, 4,5

lizzie777:

thank you so much

lizzie777:

do you want to be friends

darkknight:

No problem

darkknight:

Ya got any other questions?

lizzie777:

I''l let you know when I get back